Parallel light consisting of 2 monochromatic radiations of wavelengths 6.0 x 10^(-5) cm and 4.0 x 10^(-5) cm falls normally on a plane transmission grating ruled with 5000 lines per cm. What is the angular separation of the second-order spectra of the two wavelengths?


1
Expert's answer
2021-03-05T07:27:44-0500

Given,

Wavelengths are "(\lambda_1)=6.0\times10^{-5} cm"


"\lambda_2=4.0\times10^{-5} cm"

Grating rule "\frac{n}{d}=5000" lines per cm


"\Rightarrow \sin\theta= \frac{n\Delta \lambda}{d}"


"\Rightarrow \sin\theta = 5000\times \lambda_2-\lambda_1"

"=5000\times(6\times 10^{-5}-4\times 10^{-5})"

"=5000\times 2\times 10^{-5}"

"=0.5"

"\sin\theta = \frac{1}{2}=\sin 30^\circ"


Hence, "\theta = 30^\circ"

Hence, the required angular separation be "30^\circ"